OpenStudy (ny,ny):

quadratic functions: after finding the axis of symmetry, vertex, and y-intercept, how do you find other points on the parabola?

OpenStudy (anonymous):

Plug a value in for x to get your y value Example: f(x) = (X-2)^2 - 4 f(1) = (1-2)^2 -4 f(-1) = (-1-2)^2-4 ect.

OpenStudy (anonymous):



OpenStudy (ny,ny):

any value can work for x?

OpenStudy (anonymous):

yes

OpenStudy (ny,ny):

thanks.

OpenStudy (anonymous):

np hope it helped.

OpenStudy (ny,ny):

it did. :)